Cost of Common Equity

The future earnings, dividends, and common stock price of Callahan Technologies Inc. are expected to grow 4% per year. Callahan's common stock currently sells for $21.00 per share; its last dividend was $2.50; and it will pay a $2.60 dividend at the end of the current year.

  1. Using the DCF approach, what is its cost of common equity? Round your answer to two decimal places. Do not round your intermediate calculations.
    %
  2. If the firm's beta is 2.00, the risk-free rate is 5%, and the average return on the market is 14%, what will be the firm's cost of common equity using the CAPM approach? Round your answer to two decimal places.
    %
  3. If the firm's bonds earn a return of 9%, based on the bond-yield-plus-risk-premium approach, what will be rs? Use the midpoint of the risk premium range discussed in Section 10-5 in your calculations. Round your answer to two decimal places.
    %
  4. If you have equal confidence in the inputs used for the three approaches, what is your estimate of Callahan's cost of common equity? Round your answer to two decimal places. Do not round your intermediate calculations.
    %
Answer #1

a, Cost of equity using DCF = Dividend Next Year/Price + Growth = 2.60/21 + 4% = 16.381% or 16.38%

b. CAPM method Cost of Equity = Risk Free rate + Beta*(Market Return - Risk Free Rate) = 5%+2*(14%-5%) = 23%

c. As per the book the midpoint is 4% (Range is 3%-5%)
Bond Yield + Risk Premium = 9% +4% = 13%

d. Equal confidence in the inputs of three approaches = (16.381%+23%+13%)/3 =17.46%